# MARVL: Multi-Stage Guidance for Robotic Manipulation via Vision-Language Models

### v1

@@ 1 Introduction
Reinforcement learning (RL) has achieved notable progress in robotic manipulation (Levine et al., 2016; Kalashnikov et al., 2018; Gu et al., 2017; Han et al., 2023; Kroemer et al., 2021; Tang et al., 2025; Singh et al., 2022), yet reward specification remains a bottleneck (Yu et al., 2025; Ibrahim et al., 2024; Fu et al., 2024a; Kaelbling et al., 1996; Li, 2017; Eschmann, 2021; Lee et al., 2026). Vision-Language Models (VLMs) offer a promising path toward reward design by grounding rewards in natural language (Ma et al., 2024; Rocamonde et al., 2023; Venuto et al., 2024; Lee et al., 2026; Sontakke et al., 2023; Wang et al., 2024; Venkataraman et al., 2024). A widely adopted approach computes the cosine similarity between visual embeddings of environment observations and text embeddings of task instructions, using this score as a dense reward signal (Rocamonde et al., 2023; Adeniji et al., 2023; Baumli et al., 2023; Ma et al., 2023; Fu et al., 2024b). This strategy leverages the rich knowledge of pre-trained VLMs, but its effectiveness in guiding RL remains limited.
@@
Despite the advancements introduced in these varied studies (Rocamonde et al., 2023; Adeniji et al., 2023; Baumli et al., 2023; Ma et al., 2023; Fu et al., 2024b; Huang et al., 2024), we find that VLM-based rewards continue to face three critical challenges: (i) Weak Spatial Grounding: VLMs struggle with cross-view consistency and are highly sensitive to camera perspectives, leading to misaligned rewards. (ii) Lack of Progress Awareness: embeddings are easily distracted by irrelevant visual factors, failing to reflect incremental progress and exposing agents to reward hacking. (iii) Semantic Misalignment: VLMs frequently misinterpret task instructions, leading to unreliable reward signals. These limitations suggest that current progress in VLM-guided RL may be less robust than initially believed.
@@ 2.1 Markov Decision Process and RL
Reinforcement learning (RL) provides a principled framework for sequential decision making under uncertainty, and is typically formalized as a Markov Decision Process (MDP) (Puterman, 2014; Sutton & Barto, 2018; Li, 2017). An MDP is defined by the tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $P(s^{\prime}|s,a)$ the transition dynamics, $r(s,a)$ the reward function, and $\gamma\in[0,1)$ a discount factor. At each timestep $t$, the agent observes a state $s_{t}\in\mathcal{S}$, selects an action $a_{t}\in\mathcal{A}$, and transitions to the next state $s_{t+1}$ according to $P$, while receiving a scalar reward $r(s_{t},a_{t})$. The objective of RL is to learn a policy $\pi(a|s)$ that maximizes the expected discounted return $\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})]$. In practice, this policy is typically modelled using a neural network $\pi_{\theta}(a|s)$ with learnable parameters $\theta$ (Mnih et al., 2013). This formulation has proven effective in domains ranging from robotics to game playing, yet the design of the reward function $r$ remains a central challenge, particularly in tasks where dense supervision is difficult to obtain (Yu et al., 2025; Eschmann, 2021; Dewey, 2014).
@@ 2.2 Vision-Language Models (VLMs) and RL
Recent progress in vision-language models (VLMs) has demonstrated their ability to bridge vision and natural language through joint embedding spaces (Radford et al., 2021; Alayrac et al., 2022; Jia et al., 2021; Li et al., 2022). A landmark example is CLIP (Radford et al., 2021), which learns aligned representations of texts and images, enabling strong zero-shot transfer across diverse vision tasks. While such models were originally developed with general vision applications in mind, growing attention has been directed toward their integration with reinforcement learning (RL). VLMs provide rich semantic grounding that can benefit RL in multiple ways, including reward specification and shaping (Mahmoudieh et al., 2022; Rocamonde et al., 2023; Adeniji et al., 2023; Baumli et al., 2023; Lubana et al., 2023; Dang et al., 2023; Huang et al., 2024; Fu et al., 2024b), task description and success verification (Du et al., 2023), and state representation learning (Chen et al., 2024). These directions highlight the potential of VLMs as powerful priors that enhance vision-based RL, offering richer representations and more flexible reward signals for complex tasks.
@@ 2.3 VLMs as Rewards
Leveraging Vision-Language Models (VLMs) as a source of reward has emerged as a rapidly growing direction in reinforcement learning (Mahmoudieh et al., 2022; Rocamonde et al., 2023; Adeniji et al., 2023; Baumli et al., 2023; Lubana et al., 2023; Dang et al., 2023; Nam et al., 2023; Fu et al., 2024b; Venuto et al., 2024; Lee et al., 2026; Sontakke et al., 2023; Wang et al., 2024). Some approaches directly treat VLM-based similarity scores between observations and task instructions as the primary reward signal (Mahmoudieh et al., 2022; Rocamonde et al., 2023; Adeniji et al., 2023), while others use VLM-derived rewards as an additional source of supervision to complement the environment’s original reward (Rocamonde et al., 2023; Baumli et al., 2023; Lubana et al., 2023; Dang et al., 2023; Fu et al., 2024b). Together, these studies highlight the growing interest in integrating the rich knowledge encoded in VLMs into reinforcement learning reward design.
@@
To mitigate this challenge, recent works propose augmenting the sparse task reward with an auxiliary signal derived from VLMs (Chen et al., 2024; Rocamonde et al., 2023; Fu et al., 2024b). Concretely, the reward is reshaped as:
@@ 2.4 Challenges in VLM Rewards
Despite its simplicity, the CLIP-based reward signal is often criticized for its fuzziness, as existing literature (Fu et al., 2024b; Lubana et al., 2023; Huang et al., 2024; Mahmoudieh et al., 2022; Rocamonde et al., 2023) demonstrates that such signals are frequently compromised by semantic misalignment and false positives, failing to provide the robust feedback required for complex tasks. For instance, high similarity scores can be triggered by spurious visual–text matches that are unrelated to task progress, leading the agent toward ineffective exploration. A concrete example of such noisy and uninformative reward dynamics is shown in Figure 2 left.
To address these reliability issues, recent work has explored differing strategies. One direction focuses on suppressing irrelevant noise. BIMI (Huang et al., 2024) was introduced to mitigate false positives, and VLM-RMs-GB (Rocamonde et al., 2023) proposed Goal-Baseline Regularization to project out extraneous information. Despite their contributions, BIMI has not effectively solved the false positive problem, and VLM-RMs-GB in general lacks the adaptability required for complex tasks. Alternatively, approaches like LIV (Ma et al., 2023) and FuRL (Fu et al., 2024b) attempt to fine-tune the reward representations directly through value-oriented objectives or alignment. LIV trains embeddings with a value-oriented objective, and can optionally be fine-tuned on tailored datasets for domain adaptation, while FuRL applies lightweight alignment to the reward space. However, as also illustrated in Figure 2, these approaches remain imperfect.
While these studies provide valuable insights, their analyses remain fragmented: some highlight noise and instability (Huang et al., 2024; Lubana et al., 2023; Rocamonde et al., 2023), others note misalignment (Fu et al., 2024b; Ma et al., 2023), but none offer a comprehensive perspective. In this work, we identify three fundamental challenges underlying VLM rewards—weak spatial grounding, lack of task progress sensitivity, and semantic misalignment—which we discuss systematically in this section.
@@
Lack of Progress Awareness. While minor fluctuations are expected, a dense reward signal should exhibit a strong positive correlation with task progress (Fu et al., 2024b; Ma et al., 2023; Huang et al., 2024). However, as shown in Figure 2 (left), VLM rewards along an oracle trajectory display excessive volatility, failing to capture the global trend of task completion. The signal frequently oscillates or drops even as the agent approaches the goal. This poor signal-to-noise ratio indicates that the VLM is insensitive to fine-grained physical states, rendering the reward unreliable for guiding policy learning.
Semantic Misalignment. Leveraging VLMs for reward design assumes a well-aligned latent space where instructions naturally map to their corresponding observations (Fu et al., 2024b; Huang et al., 2024; Lubana et al., 2023; Sontakke et al., 2023; Ma et al., 2023). However, Figure 2 (right) contradicts this assumption: the heatmap exhibits diffuse activations with only a weak diagonal, indicating poor task separability in the embedding space. Consistently, Figure 2 (left) shows that even after adaptations such as LIV and FuRL, the reward remains strictly negative, suggesting a persistent representation misalignment. In this setting, the reward provides only a weak ordinal signal—ranking relative similarities among poorly aligned instances—rather than a calibrated measure of semantic correctness.
@@ 4.2 Baselines
3. FuRL (Fu et al., 2024b). A recent VLM-based reward method that leverages two key mechanisms: lightweight reward alignment and relay RL, where a VLM-guided agent and an auxiliary SAC agent take turns acting in the environment and share a replay buffer so that the SAC agent can help the VLM agent escape local minima induced by fuzzy VLM rewards.

### v3

@@ 1 Introduction
Reinforcement learning (RL) has achieved notable progress in robotic manipulation (Levine et al., 2016; Kalashnikov et al., 2018; Gu et al., 2017; Han et al., 2023; Kroemer et al., 2021; Tang et al., 2025; Singh et al., 2022), yet reward specification remains a bottleneck (Yu et al., 2025; Ibrahim et al., 2024; Fu et al., 2024b; Kaelbling et al., 1996; Li, 2017; Eschmann, 2021; Lee et al., 2026). Vision-Language Models (VLMs) offer a promising path toward reward design by grounding rewards in natural language (Ma et al., 2024; Rocamonde et al., 2023; Venuto et al., 2024; Lee et al., 2026; Sontakke et al., 2023; Wang et al., 2024; Venkataraman et al., 2024). A widely adopted approach computes the cosine similarity between visual embeddings of environment observations and text embeddings of task instructions, using this score as a dense reward signal (Rocamonde et al., 2023; Adeniji et al., 2023; Baumli et al., 2023; Ma et al., 2023; Fu et al., 2024a). This strategy leverages the rich knowledge of pre-trained VLMs, but its effectiveness in guiding RL remains limited.
@@
Despite the advancements introduced in these varied studies (Rocamonde et al., 2023; Adeniji et al., 2023; Baumli et al., 2023; Ma et al., 2023; Fu et al., 2024a; Huang et al., 2024), we find that VLM-based rewards continue to face three critical challenges: (i) Weak Spatial Grounding: VLMs struggle with cross-view consistency and are highly sensitive to camera perspectives, leading to misaligned rewards. (ii) Lack of Progress Awareness: embeddings are easily distracted by irrelevant visual factors, failing to reflect incremental progress and exposing agents to reward hacking. (iii) Semantic Misalignment: VLMs frequently misinterpret task instructions, leading to unreliable reward signals. These limitations suggest that current progress in VLM-guided RL may be less robust than initially believed.
@@ 2.1 Markov Decision Process and RL
Reinforcement learning (RL) provides a principled framework for sequential decision making under uncertainty, and is typically formalized as a Markov Decision Process (MDP) (Puterman, 2014; Sutton and Barto, 2018; Li, 2017). An MDP is defined by the tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $P(s^{\prime}|s,a)$ the transition dynamics, $r(s,a)$ the reward function, and $\gamma\in[0,1)$ a discount factor. At each timestep $t$, the agent observes a state $s_{t}\in\mathcal{S}$, selects an action $a_{t}\in\mathcal{A}$, and transitions to the next state $s_{t+1}$ according to $P$, while receiving a scalar reward $r(s_{t},a_{t})$. The objective of RL is to learn a policy $\pi(a|s)$ that maximizes the expected discounted return $\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})]$. In practice, this policy is typically modelled using a neural network $\pi_{\theta}(a|s)$ with learnable parameters $\theta$ (Mnih et al., 2013). This formulation has proven effective in domains ranging from robotics to game playing, yet the design of the reward function $r$ remains a central challenge, particularly in tasks where dense supervision is difficult to obtain (Yu et al., 2025; Eschmann, 2021; Dewey, 2014).
@@ 2.2 Vision-Language Models (VLMs) and RL
Recent progress in vision-language models (VLMs) has demonstrated their ability to bridge vision and natural language through joint embedding spaces (Radford et al., 2021; Alayrac et al., 2022; Jia et al., 2021; Li et al., 2022). A landmark example is CLIP (Radford et al., 2021), which learns aligned representations of texts and images, enabling strong zero-shot transfer across diverse vision tasks. While such models were originally developed with general vision applications in mind, growing attention has been directed toward their integration with reinforcement learning (RL). VLMs provide rich semantic grounding that can benefit RL in multiple ways, including reward specification and shaping (Mahmoudieh et al., 2022; Rocamonde et al., 2023; Adeniji et al., 2023; Baumli et al., 2023; Lubana et al., 2023; Dang et al., 2023; Huang et al., 2024; Fu et al., 2024a), task description and success verification (Du et al., 2023), and state representation learning (Chen et al., 2024). These directions highlight the potential of VLMs as powerful priors that enhance vision-based RL, offering richer representations and more flexible reward signals for complex tasks.
@@ 2.3 VLMs as Rewards
Leveraging Vision-Language Models (VLMs) as a source of reward has emerged as a rapidly growing direction in reinforcement learning (Mahmoudieh et al., 2022; Rocamonde et al., 2023; Adeniji et al., 2023; Baumli et al., 2023; Lubana et al., 2023; Dang et al., 2023; Nam et al., 2023; Fu et al., 2024a; Venuto et al., 2024; Lee et al., 2026; Sontakke et al., 2023; Wang et al., 2024). Some approaches directly treat VLM-based similarity scores between observations and task instructions as the primary reward signal (Mahmoudieh et al., 2022; Rocamonde et al., 2023; Adeniji et al., 2023), while others use VLM-derived rewards as an additional source of supervision to complement the environment’s original reward (Rocamonde et al., 2023; Baumli et al., 2023; Lubana et al., 2023; Dang et al., 2023; Fu et al., 2024a). Together, these studies highlight the growing interest in integrating the rich knowledge encoded in VLMs into reinforcement learning reward design.
@@
To mitigate this challenge, recent works propose augmenting the sparse task reward with an auxiliary signal derived from VLMs (Chen et al., 2024; Rocamonde et al., 2023; Fu et al., 2024a). Concretely, the reward is reshaped as:
@@ 2.4 Challenges in VLM Rewards
Despite its simplicity, the CLIP-based reward signal is often criticized for its fuzziness, as existing literature (Fu et al., 2024a; Lubana et al., 2023; Huang et al., 2024; Mahmoudieh et al., 2022; Rocamonde et al., 2023) demonstrates that such signals are frequently compromised by semantic misalignment and false positives, failing to provide the robust feedback required for complex tasks. For instance, high similarity scores can be triggered by spurious visual–text matches that are unrelated to task progress, leading the agent toward ineffective exploration. A concrete example of such noisy and uninformative reward dynamics is shown in Figure 2 left.
To address these reliability issues, recent work has explored differing strategies. One direction focuses on suppressing irrelevant noise. BIMI (Huang et al., 2024) was introduced to mitigate false positives, and VLM-RMs-GB (Rocamonde et al., 2023) proposed Goal-Baseline Regularization to project out extraneous information. Despite their contributions, BIMI has not effectively solved the false positive problem, and VLM-RMs-GB in general lacks the adaptability required for complex tasks. Alternatively, approaches like LIV (Ma et al., 2023) and FuRL (Fu et al., 2024a) attempt to fine-tune the reward representations directly through value-oriented objectives or alignment. LIV trains embeddings with a value-oriented objective, and can optionally be fine-tuned on tailored datasets for domain adaptation, while FuRL applies lightweight alignment to the reward space. However, as also illustrated in Figure 2, these approaches remain imperfect.
While these studies provide valuable insights, their analyses remain fragmented: some highlight noise and instability (Huang et al., 2024; Lubana et al., 2023; Rocamonde et al., 2023), others note misalignment (Fu et al., 2024a; Ma et al., 2023), but none offer a comprehensive perspective. In this work, we identify three fundamental challenges underlying VLM rewards—weak spatial grounding, lack of task progress sensitivity, and semantic misalignment—which we discuss systematically in this section.
@@
Lack of Progress Awareness. While minor fluctuations are expected, a dense reward signal should exhibit a strong positive correlation with task progress (Fu et al., 2024a; Ma et al., 2023; Huang et al., 2024). However, as shown in Figure 2 (left), VLM rewards along an oracle trajectory display excessive volatility, failing to capture the global trend of task completion. The signal frequently oscillates or drops even as the agent approaches the goal. This poor signal-to-noise ratio indicates that the VLM is insensitive to fine-grained physical states, rendering the reward unreliable for guiding policy learning.
Semantic Misalignment. Leveraging VLMs for reward design assumes a well-aligned latent space where instructions naturally map to their corresponding observations (Fu et al., 2024a; Huang et al., 2024; Lubana et al., 2023; Sontakke et al., 2023; Ma et al., 2023). However, Figure 2 (right) contradicts this assumption: the heatmap exhibits diffuse activations with only a weak diagonal, indicating poor task separability in the embedding space. Consistently, Figure 2 (left) shows that even after adaptations such as LIV and FuRL, the reward remains strictly negative, suggesting a persistent representation misalignment. In this setting, the reward provides only a weak ordinal signal—ranking relative similarities among poorly aligned instances—rather than a calibrated measure of semantic correctness.
@@ 4.2 Baselines
3. FuRL (Fu et al., 2024a). A recent VLM-based reward method that leverages two key mechanisms: lightweight reward alignment and relay RL, where a VLM-guided agent and an auxiliary SAC agent take turns acting in the environment and share a replay buffer so that the SAC agent can help the VLM agent escape local minima induced by fuzzy VLM rewards.